OpenStudy (anonymous):

please help The sum of the measures of the angles of any triangle is 180 degrees. In triangle ABC, angles A and B have the same measure, while the measure of angle C is 90 degrees larger than each of A and B. What are the measures of the three angles? Angle a is Angle b is Angle C is

OpenStudy (anonymous):

\[180=2x+(x+90)\] x represents the value you don't know <A & <B. <C is 90 degrees bigger than <A or<B. A &B are your 2x and C is x+90 Solve for x

OpenStudy (anonymous):

i don't understand

OpenStudy (anonymous):

 C is drawn obtuse because it is more than 90 degrees

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

A and B are =

OpenStudy (anonymous):

Let's call the measure that we do not know x. A=x and B=x

OpenStudy (anonymous):

C=x+90 because it says "C is 90 degrees larger than each of A and B"

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

You total sum of the angles inside a triangle is 180degrees

OpenStudy (anonymous):

SO let's make an equation to find what the measure of x is

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

\[180=2x+(x+90)\]

OpenStudy (anonymous):

2x is <A + <B and (x+90) is <C

OpenStudy (anonymous):

Just solve for x

OpenStudy (anonymous):

how

OpenStudy (anonymous):

HOW? Have you never solved for x?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Then why did you this problem?

OpenStudy (anonymous):

You have to know basic algebra to do geometry

OpenStudy (anonymous):

im practicing because i am getting ready to start school

OpenStudy (anonymous):

What grade?

OpenStudy (anonymous):

10 grade

OpenStudy (anonymous):

i am not good at math so i am just practicing

OpenStudy (anonymous):

Have you ever done any problems with variables?

OpenStudy (anonymous):

no this is the first time

OpenStudy (anonymous):

What school system are you in? You should have been doing math with variables since at least 7th grade

OpenStudy (anonymous):

I live in jersey

OpenStudy (anonymous):

i have just never been good at math

OpenStudy (anonymous):

Jesus. Hon I am so sorry for you. They really have you guys behind. Okay back to our equation.

OpenStudy (anonymous):

\[180=2x+(x+90)\]

OpenStudy (anonymous):

i am trying to work on it before school starts so i do better

OpenStudy (anonymous):

with variables, the same letter to the same power are like terms

OpenStudy (anonymous):

So you can combine 2x and x into 3x, because they're both positive. so now we have \[180=3x+90\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Now you subtract 90 from both sides because what we do to one side we do to the other

OpenStudy (anonymous):

That cancels out your 90 on the right side. Now divide both sides by 3.

OpenStudy (anonymous):

The steps we are taking to get x alone is called inverse operations

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

what do you have now?

OpenStudy (anonymous):

im still trying to figure it out

OpenStudy (anonymous):

60 is that right ?

OpenStudy (anonymous):

180-90=90

OpenStudy (anonymous):

then you divide 90 by 3

OpenStudy (anonymous):

oh i divided the wrong number okay

OpenStudy (anonymous):

30

OpenStudy (anonymous):

\[180=2x+(x+90)\]combine like terms \[180=3x+90\]subtract 90 from both sides\[90=3x\]divide both sides by 3\[30=x\]

OpenStudy (anonymous):

SO <A=30, <B=30 and <C=30+90

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

okay what do you do next ?

OpenStudy (anonymous):

Those are your answers. :) <A=30 <B=30 and <C=120

OpenStudy (anonymous):

okay i want to write all this down so i can study